# Yuxin Feng4, Long Ma2†, Bo Du5, Pheng-Ann Heng1

### v1

@@ Abstract
In laparoscopic surgery, a clear and high-quality visual field is critical for surgeons to make accurate intraoperative decisions. However, persistent visual degradation, including smoke generated by energy devices, lens fogging from thermal gradients, and lens contamination due to blood or tissue fluid splashes during surgical procedures, severely impair visual clarity. These degenerations can seriously hinder surgical workflow and pose risks to patient safety. To systematically investigate and address various forms of surgical scene degradation, we introduce a real-world open-source surgical image restoration dataset covering laparoscopic environments, called SurgClean, which involves multi-type image restoration tasks, e.g., desmoking, defogging, and desplashing. SurgClean comprises 1,020 images with diverse degradation types and corresponding paired reference labels. Based on SurgClean, we establish a standardized evaluation benchmark and provide performance for 22 representative generic task-specific image restoration approaches, including 12 generic and 10 task-specific image restoration approaches. Experimental results reveal substantial performance gaps relative to clinical requirements, highlighting a critical opportunity for algorithm advancements in intelligent surgical restoration. Furthermore, we explore the degradation discrepancies between surgical and natural scenes from structural perception and semantic understanding perspectives, providing fundamental insights for domain-specific image restoration research. Our work aims to empower the capabilities of restoration algorithms to increase surgical environments and improve the efficiency of clinical procedures.
@@ 1 Introduction
Backgrounds: In clinical minimally invasive surgery, a clear surgical field of view and improved image quality in the operative area are critical for preventing unexpected complications and ensuring surgical safety. Especially in narrow laparoscopic environments, visual clarity could be severely impaired by various intraoperative factors, e.g., smoke generated by energy instruments, fog formation on lenses due to temperature differences, and lens contamination caused by splashes of blood or tissue fluids during surgical manipulations, as illustrated in Fig. 1. Diverse types of degradation seriously hinder surgeons’ observation and judgment of surgical conditions, making it difficult to accurately identify anatomy and lesion sites, potentially leading to incorrect operations and reducing the success rate of surgery. Thus, laparoscopic surgical image restoration (LSIR) is critical to providing surgeons with clear and stable intraoperative views, significantly reducing surgery time, minimizing surgical errors, and improving operational efficiency.
@@
Challenges: Unlike natural scenarios, the laparoscopic environment poses unique challenges for image restoration, characterized by diverse types and levels of image degradation, which seriously distort the surgical field of view and significantly hinder the distinction of anatomical structures [11, 23]. As illustrated in Fig. 1, Smoke originates from numerous fine tissue fragments and fat vaporization when surgeons use energy devices to cauterize diseased areas and surrounding tissues. Fog is mainly caused by condensation of water vapor when the temperature difference between the inside and outside of the abdominal cavity leads to mist forming on the lens. Compared to smoke noise, the distribution of surgical fog is broader and more uniform. For Splashing, the use of instruments such as electrocautery and ultrasonic scalpels during surgery results in the splashing of blood, tissue fluid, and other substances onto the lens, causing contamination and increasing the difficulty of restoration. Moreover, the restoration process requires cleaning anatomical details, e.g., blood vessels and tissue structures, to assist doctors in making precise surgical decisions. When complex degradations are superimposed, critical clues obscured by noisy substances may blend into the background, further increasing degradation complexity. To this end, LSIR presents unique complexities and challenges, highlighting the urgency of developing dedicated benchmark datasets for surgical restoration.
Figure 1: Illustration of surgical image restoration in laparoscopic surgery, including laparoscopic operating scenes, multi-type degradation causes and images, and algorithms for surgical restoration.
@@
Datasets: Existing LSIR datasets usually have limitations in terms of authenticity and diversity. Several surgical restoration datasets [25, 4, 30] are generated through manual synthesis, resulting in discrepancies with real-world data and restricting the applicability of algorithms in actual scenarios. Other datasets derived from real surgical scenes [28, 21, 31] mainly focus on a single type of degradation, such as desmoking. To comprehensively investigate degradation challenges in real-world laparoscopic surgery, we construct the first multi-source LSIR dataset, SurgClean, aimed at effectively assisting surgeons in reducing visual noise in the operating field of view. SurgClean comprises 1,020 laparoscopic images under various degradation conditions selected from real surgical videos of 134 patients, accompanied by adjacent clean frames as paired reference labels. Driven by clinical needs, our dataset introduces three challenging LSIR tasks: desmoking, defogging, and desplashing, and further subdivides the degradation levels according to difficulty and characteristics. This dataset enables researchers to systematically explore and evaluate various image restoration methods under real laparoscopic conditions, providing opportunities for broader research and algorithm development.
@@
Benchmarks: We establish a comprehensive LSIR benchmark to address the inherent data, properties, and computational challenges based on SurgClean. For desmoking, defogging, and desplashing, we evaluate 22 representative image restoration methods, including general and task-specific frameworks. Extensive experimental results reveal that: 1) Although existing image restoration techniques perform effectively in natural scenes, their accuracy and efficiency are insufficient to meet clinical requirements when dealing with complex multi-factor degradation. 2) While current methods show reasonable performance for samples with lower-level degradation, they fail to produce noticeable improvements for severe degradation at higher levels. These findings highlight the significant potential and necessity for developing advanced algorithms specifically designed for LSIR. By contributing to this benchmark community, we aim to guide future research prospects and methodological innovations, ultimately enhancing the efficiency of minimally invasive surgery and reducing surgical risks.
@@
The main contributions of this study are three-fold: (1) We conducted a thorough exploration of various types of degradation in real-world surgical scenarios and introduced SurgClean, the first multi-source LSIR dataset in laparoscopic surgery. (2) We established a comprehensive LSIR benchmark based on the SurgClean dataset, achieving a performance evaluation of multiple types and levels of image restoration methods. (3) We revisited various degradation characteristics specific to surgical conditions and performed a manifold discrepancy analysis with degradation in natural scenes, intending to explore potential optimization strategies for surgical degradation.
@@ 2 Related Works
Surgical Restoration Datasets. Various surgical image restoration (SIR) datasets [25, 4, 30, 28, 21, 33] have been developed to support the training and evaluation of restoration models in surgical scenarios. Among them, datasets such as [25, 4, 3, 30] are generated via manual synthesis during post-processing, leading to notable discrepancies from real-world surgical degradations and limiting their applicability to clinical practice. For instance, Chen et al. [4] proposed the SmokeGCN dataset, which applies Gaussian filters to generate various smoke effects and overlay them on clean images. While synthetic datasets have facilitated early research efforts in SIR, they struggle to adapt to the complexity and variability of clinical environments, thereby limiting the generalization ability of models. In contrast, datasets like Cyclic-DesmokeGAN [28], Desmoke-LAP [21], DesmokeData [33], and LSVD [31] are derived from real endoscopic surgeries and focus on addressing desmoking task. DeSmoke-LAP [21] focuses on desmoking in hysterectomy videos using an unpaired image-to-image translation approach. LSVD [31] comprises various smoke scenes from real laparoscopic surgery videos and adopts pre-smoke frames as supervision references. Although these datasets provide valuable resources for specific restoration tasks, they mainly focus on isolated scenarios and lack diversity in degradation types. To bridge this gap, we present a new comprehensive dataset to systematically explore multiple real-world surgical degradation challenges.
Image Restoration Approaches. General image restoration methods can be categorized into three categories: Convolutional Neural Network (CNN)-based, Transformer-based, and Mamba-based architectures. CNN-based methodologies like ConvIR [9] and FocalNet [8] demonstrate robust performance across diverse restoration tasks by utilizing multi-scale feature extraction capabilities. Transformer-based techniques such as Restormer [36], X-Restormer [5], AST [40], RAMiT [7], and Histoformer [27] improves restoration quality through attention mechanisms that enhance their modeling capacity. Recent advancements in state space models, particularly with the emergence of the Mamba architecture, have propelled Mamba-based modeling into prominence across various vision tasks including image restoration, with notable implementations like MambaIR [12] and MambaIRv2 [13]. As for the defogging task: RIDCP [32] proposes an enhanced haze imaging model integrated with codebook-based methodology to achieve robust image dehazing. C2PNet [38] implements contrastive learning to strengthen discriminative capability. DEA [6] focuses on designing a more effective network architecture to extract robust features. CORUN [10] develops a collaborative unfolding network, while CoA [20] employs a compress-then-adapt strategy to bridge synthetic-to-real domain gaps. For the desmoking task: DCP[14] serves as a classical approach that offers a simple yet effective solution for smoke removal. MS-CNN [29] integrates Laplacian pyramid decomposition with CNN architecture to perform desmoking in medical imaging scenarios. IIT-EDC [24] utilizes a lightweight GAN framework for real-time processing, while Desmoke-LAP[21] builds upon CycleGAN for medical smoke elimination. SelfSVD [31] implements self-supervision by utilizing clean frames from smoke-contaminated videos to guide the desmoking process.
@@ 3 SurgClean Dataset
Stemming from clinical requirements, our SurgClean dataset provides three challenging laparoscopic surgical image restoration (LSIR) tasks in real-world clinical scenarios, i.e., desmoking, defogging, and desplashing. We provide a detailed overview of SurgClean in terms of data collection, data annotation, and statistics analysis.
@@ 3.1 Data Collection
The degradation of surgical scenes caused by contaminants such as tissue fragments and smoke during laparoscopy occurs according to the specific surgical procedure and has a certain degree of randomness. Therefore, we focus only on restoring noise images that appear at specific time points rather than video-level restoration to improve clinical application efficiency. Concretely, we collect laparoscopic surgical videos of 134 patients undergoing hepatectomy and cholecystectomy surgeries from partner hospitals to construct the SurgClean dataset. Four intern surgeons were invited to pick out the frames from each video where they identified the surgical field as obscured by visual interference, including smoke, fog, and splash contamination on the lens. Meanwhile, two professional surgeons review and filter all interfering image frames to ensure the reliability of the samples. Ultimately, we gather a total of 1,020 noisy images with a resolution of 1280$\times$720, including 600 images for desmoking, 320 images for defogging, and 100 images for the desplashing task. The data in each sub-task is divided into training and test sets in an 8:2 ratio. Fig. 3 illustrates the data distribution across the three types of LSIR in SurgClean. The proportion of samples taken up by different sub-tasks also reflects the real-world occurrence rates of interference events during hepatectomy and cholecystectomy procedures. For instance, the probability of interference due to splashing is the lowest among the three types. Fig. 3 also exhibits various complex examples of visual interference encountered in laparoscopic surgery.
Figure 3: Illustration of the proposed SurgClean dataset. We exhibit data distribution under different degradation scenarios and further provide fine-grained divisions according to interference levels and attributes. Typical samples are displayed on both sides, zoom in for details.
@@ 3.2 Data Annotations
Since all samples in SurgClean are derived from real-world laparoscopic surgical videos, it is difficult to obtain perfectly aligned ground truth frames. To provide relatively clear paired reference labels for deep learning-based model supervision, we adopt the PS-frame scheme in line with [31], i.e., using the last clean frame preceding the noisy frame as the paired reference. As shown in Fig. 2, this form of label pair may be affected by the movement of the laparoscopic lens, resulting in slight deviations in pixel distribution. This also poses a key challenge for model supervision in this benchmark. To reduce the impact of misaligned references on the restoration model during training, we uniformly employ the optical flow alignment method [37, 26] to estimate the optical flow between the input image and the unaligned reference for relative displacement. Then, we warp the reference using the estimated optical flow to transform the unaligned images into aligned pairs. This strategy allows us to train and evaluate models as same as general image restoration tasks for fairness and effectiveness.
@@ 3.3 Data Statistics and Analysis
Table 1: Degradation level description.
Level Descriptions
Level 1
Mild degradation: Less than 1/3 of the entire field of view is filled
with smoke or fog, and the anatomical structures in the operating area
are visible and do not affect judgment.
Level 2
Moderate degradation: 1/3-2/3 of the entire field of view is filled with
smoke or fog, and the anatomical structures in the operating area are
still visible but less clearly than in Level 1, without affecting judgment.
Level 3
Severe degradation: More than 2/3 of the entire field of view is filled
with smoke or fog, or the operating area
is degraded, which may affect judgment.
Level 4
Complete degradation: More than 2/3 of the entire field of view is filled
with smoke or fog, and the operating area is degraded, severely affecting
judgment. Surgery must be paused to wait for the degradation to clear.
To facilitate more fine-grained research, we further divided the dataset into multiple subsets based on different degradation levels and attributes. Specifically, for desmoking and defogging tasks, surgeons divide corresponding samples into four levels (from easy to difficult, i.e., Level 1-4) based on the area and degree of smoke or fog obstruction and the degree of interference with the operating zone. For desplashing, the samples are grouped into four subsets (i.e., $T_{blood}$, $T_{fat}$, $T_{bile}$, and $T_{fluid}$) based on the type of substances splashed onto the lens, including blood, fat, bile, and tissue fluid. We clearly elaborate on the definition and classification of different levels of degradation in Table 1. In addition, Fig. 3 exhibits typical examples of different difficulty levels and each type of splash interference. More sample exhibitions and information statistics of our SurgClean dataset can be found in Appendix. B.
Table 2: Comparison between SurgClean and existing real-world surgical image restoration datasets.‘OA’ stands for whether the data are publicly available.
Dataset OA Paired Label Resolution Desmoke Defog Desplash Scale Source No. Source
Cyclic-DesmokeGAN [28] 240$\times$320 1,400 1 Cholec80
Desmoke-LAP [21] 720$\times$540 3,000 1 Hysterectomy
DesmokeData [33] 700$\times$350 961 1 da Vinci Si videos
SurgClean (Ours) 1280$\times$720 1,020 4 Cholecystectomy, Hepatectomy, Pancreatectomy, Splenectomy
Furthermore, we summarize and perform comparisons with other real-world LSIR datasets. As illustrated in Table 2, existing LSIR datasets focus only on the desmoking task, while SurgClean comprehensively embraces multiple types of image degradation in actual surgical procedures. In addition, datasets with paired reference labels provide more significant support for model fine-tuning. Compared to other datasets with a single data source, our data source comes from various types of laparoscopic surgery, e.g., hepatectomy, cholecystectomy, and splenectomy, to accommodate data diversity. SurgClean comprises a variety of interference scenarios in laparoscopic surgery to promote comprehensive evaluation and advancement in surgical image restoration tasks. Additionally, the presence of paired reference labels facilitates the training and evaluation of deep learning-based restoration models. We conduct cross-validation experiments with DesmokeData [33] to validate the effectiveness and challenging nature of our dataset (refer to Appendix. E). Meanwhile, we also perform a visual comparison of degraded images in our dataset with other LSIR datasets in Appendix. C. Our dataset has passed the ethics statement and can be publicly available at https://github.com/PJLallen/Surgical-Image-Restoration.

### v2

# Benchmarking Endoscopic Surgical Image Restoration and Beyond

@@ Abstract
In endoscopic surgery, a clear and high-quality visual field is critical for surgeons to make accurate intraoperative decisions. However, persistent visual degradation, including smoke generated by energy devices, lens fogging from thermal gradients, and lens contamination due to blood or tissue fluid splashes during surgical procedures, severely impairs visual clarity. These degenerations can seriously hinder surgical workflow and pose risks to patient safety. To systematically investigate and address various forms of surgical scene degradation, we introduce a real-world open-source surgical image restoration dataset covering endoscopic environments, called SurgClean, which involves multi-type image restoration tasks from two medical sites, i.e., desmoking, defogging, and desplashing. SurgClean comprises 3,113 images with diverse degradation types and corresponding paired reference labels. Based on SurgClean, we establish a standardized evaluation benchmark and provide performance for 22 representative generic task-specific image restoration approaches, including 12 generic and 10 task-specific image restoration approaches. Experimental results reveal substantial performance gaps relative to clinical requirements, highlighting a critical opportunity for algorithm advancements in intelligent surgical restoration. Furthermore, we explore the degradation discrepancies between surgical and natural scenes from structural perception and semantic understanding perspectives, providing fundamental insights for domain-specific image restoration research. Our work aims to empower restoration algorithms and improve the efficiency of clinical procedures. Data and code are available.
@@ 1 Introduction
Backgrounds: In minimally invasive surgery, a clear surgical field of view and improved image quality in the operative area are critical for preventing unexpected complications and ensuring surgical safety. Especially in narrow endoscopic environments, visual clarity could be severely impaired by various intraoperative factors, e.g., smoke generated by energy instruments, fog formation on lenses due to temperature differences, and lens contamination caused by splashes of blood or tissue fluids during surgical manipulations, as illustrated in Fig. LABEL:fig:fig1. Diverse types of degradation seriously hinder surgeons’ observation and judgment of surgical conditions, making it difficult to accurately identify anatomy and lesion sites, potentially leading to incorrect operations and reducing the success rate of surgery. Thus, endoscopic surgical image restoration (ESIR) is critical to providing surgeons with clear and stable intraoperative views, significantly reducing surgery time, minimizing surgical errors, and improving operational efficiency.
@@
Challenges: Unlike natural scenarios, the endoscopic environment poses unique challenges for image restoration, characterized by diverse types and levels of image degradation, which seriously distort the surgical field of view and significantly hinder the distinction of anatomical structures [guo2025surgical, pei2024depth]. As illustrated in Fig. LABEL:fig:fig1, Smoke originates from numerous fine tissue fragments and fat vaporization when surgeons use energy devices to cauterize diseased areas and surrounding tissues. Fog is mainly caused by condensation of water vapor when the temperature gradient between the intra- and extra-corporeal environments leads to mist forming on the lens. Compared to smoke noise, the distribution of surgical fog is broader and more uniform. For Splashing, the use of instruments such as electrocautery and ultrasonic scalpels during surgery results in the splashing of blood, tissue fluid, and other substances onto the lens, causing contamination and increasing the difficulty of restoration. Moreover, the restoration process requires cleaning anatomical details, e.g., blood vessels and tissue structures, to assist doctors in making precise surgical decisions. When complex degradations are superimposed, critical clues obscured by noisy substances may blend into the background, further increasing degradation complexity. To this end, ESIR presents unique complexities and challenges, highlighting the urgency of developing dedicated benchmark datasets for surgical restoration.
@@
Datasets: Existing ESIR datasets usually have limitations in terms of authenticity and diversity. Several surgical restoration datasets [sengar2021multi, chen2019smokegcn, wang2023surgical] are generated through manual synthesis, restricting the applicability of algorithms in actual scenarios. Other datasets derived from real surgical scenes [venkatesh2020unsupervised, pan2022desmoke, wu2025self] mainly focus on a single type of degradation, such as desmoking. To comprehensively investigate degradation challenges in real-world endoscopic surgery, we construct the first multi-source ESIR dataset, SurgClean, aimed at effectively assisting surgeons in reducing visual noise in the operating field of view. SurgClean comprises 3,113 endoscopic images under various degradations selected from real surgical videos of 414 patients across two medical sites, accompanied by adjacent clean frames as paired reference labels. Driven by clinical needs, our dataset introduces three challenging ESIR tasks: desmoking, defogging, and desplashing, and further subdivides the degradation levels according to difficulty and characteristics. This dataset enables researchers to systematically explore and evaluate various image restoration methods under real endoscopic conditions, providing opportunities for broader research and algorithm development.
@@
Benchmarks: We establish a comprehensive ESIR benchmark to address the inherent data, properties, and computational challenges based on SurgClean. For desmoking, defogging, and desplashing, we evaluate 22 representative image restoration methods, including general and task-specific frameworks. Extensive experimental results reveal that: 1) Although existing image restoration techniques perform effectively in natural scenes, their accuracy and efficiency are insufficient to meet clinical requirements when dealing with complex multi-factor degradation. 2) While current methods show reasonable performance for samples with lower-level degradation, they fail to produce noticeable improvements for severe degradation at higher levels. These findings highlight the significant potential and necessity for developing advanced algorithms specifically designed for ESIR. By contributing to this benchmark community, we aim to guide future research prospects and methodological innovations, ultimately enhancing the efficiency of minimally invasive surgery and reducing surgical risks.
@@
Our main contributions are three-fold: (1) We conducted a thorough exploration of various types of degradation in real-world surgical scenarios and introduced SurgClean, the first multi-source ESIR dataset in endoscopic surgery. (2) We established a comprehensive ESIR benchmark based on SurgClean, achieving a performance evaluation of multiple types and levels of restoration methods. (3) We revisited various degradation characteristics specific to surgical conditions and performed a manifold discrepancy analysis with degradation in natural scenes, intending to explore potential optimization strategies for surgical degradation.
@@ 2 Related Works
Figure 3: Illustration of the proposed SurgClean dataset. We exhibit data distribution under different degradation scenarios and further provide fine-grained divisions according to interference levels and attributes. Typical samples are displayed on both sides.
Surgical Restoration Datasets. Numerous surgical image restoration (SIR) datasets have been introduced for training and evaluation in surgical scenes [sengar2021multi, chen2019smokegcn, wang2023surgical, venkatesh2020unsupervised, pan2022desmoke, xia2024new]. Many are synthetic post-processed sets [sengar2021multi, chen2019smokegcn, chang2024lsd3k, wang2023surgical], creating a gap to clinical degradations; for example, SmokeGCN [chen2019smokegcn] overlays Gaussian-filtered smoke on clean images. These datasets enabled early progress but generalize poorly to complex, variable clinical conditions. More recent real-surgery datasets emphasize desmoking, including Cyclic-DesmokeGAN [venkatesh2020unsupervised], DeSmoke-LAP [pan2022desmoke], DesmokeData [xia2024new], and LSVD [wu2025self]; DeSmoke-LAP targets hysterectomy videos with unpaired translation, and LSVD mines smoke scenes using pre-smoke frames as references. Despite their value, these resources typically address a single degradation and offer limited diversity. To bridge this gap, we introduce a comprehensive dataset covering multiple real-world surgical degradations.
Image Restoration Approaches. General image restoration methods broadly fall into three families: CNN-based, Transformer-based, and Mamba-based architectures. CNN models (e.g., ConvIR [cui2024revitalizing], FocalNet [focal2023]) leverage multi-scale feature extraction for robust restoration. Transformer variants (Restormer [zamir2022restormer], X-Restormer [chen2024xrestormer], AST [zhou2024AST], RAMiT [choi2024reciprocal], Histoformer [sun2025restoring]) enhance modeling capacity via attention. Recent state-space models, notably Mamba, have motivated Mamba-based restorers such as MambaIR [guo2025mambair] and MambaIRv2 [guo2024mambairv2]. For defogging, methods include RIDCP [wu2023ridcp] with an enhanced haze model and codebook design, C2PNet [zheng2023curricular] with contrastive learning, DEA [chen2024dea] with stronger feature extractors, CORUN [fang2024real] with collaborative unfolding, and CoA [ma2025coa] with a compress-then-adapt pipeline for synthetic-to-real transfer. For desmoking, DCP [he2010single] remains a simple baseline; MS-CNN [wang2019multiscale] combines Laplacian pyramids with CNNs; IIT-EDC [salazar2020desmoking] targets real-time GAN-based removal; Desmoke-LAP [pan2022desmoke] builds on CycleGAN; SelfSVD [wu2025self] implements self-supervision by utilizing clean frames from smoke-contaminated videos to guide the desmoking process. However, a unified restoration framework that robustly handles multiple surgical degradations remains largely unexplored.
@@ 3 SurgClean Dataset
Table 1: Comparing SurgClean with existing real-world surgical image restoration datasets. “OA” denotes public availability.
Dataset OA Reference Resolution Desmoke Defog Desplash Scale Site No. Data Source
[…]
  • Site A: {Gall: Gallbladder, Bile: Bile Duct, Pan: Pancreatectomy, Spl: Splenectomy, Liver}; Site B: {Med: Mediastinal, Eso: Esophageal, Lung}.
@@ 3.1 Data Collection
[…]
@@ 3.2 Data Annotations
Since all samples in SurgClean are derived from real-world endoscopic surgical videos, it is difficult to obtain perfectly aligned ground truth frames. To provide relatively clear paired reference labels for deep learning-based model supervision, we adopt the PS-frame scheme in line with [wu2025self], i.e., using the last clean frame preceding the noisy frame as the paired reference. As shown in Fig. 2, this form of pairing may be affected by the movement of the endoscopic lens, resulting in slight deviations in pixel distribution. This also poses a key challenge for model supervision in this benchmark. To reduce the impact of misaligned references on the restoration model during training, we uniformly employ the optical flow alignment method [zhang2021learning, sun2018pwc] to estimate the optical flow between the input image and the unaligned reference for relative displacement. Then, we warp the reference using the estimated optical flow to transform the unaligned images into aligned pairs. This strategy allows us to train and evaluate models in as same as general image restoration tasks for fairness and effectiveness.
Table 2: Degradation level description in SurgClean.
Level Detailed Descriptions
Level 1 Mild degradation: Less than 1/3 of the entire field of view is filled with smoke or fog, and the anatomical structures in the operating area are visible and do not affect judgment.
Level 2 Moderate degradation: 1/3–2/3 of the entire field of view is filled with smoke or fog, and the anatomical structures in the operating area are still visible but less clearly than in Level 1, without affecting judgment.
Level 3 Severe degradation: More than 2/3 of the entire field of view is filled with smoke or fog, or the operating area is degraded, which may affect judgment.
Level 4 Complete degradation: More than 2/3 of the entire field of view is filled with smoke or fog, and the operating area is degraded, severely affecting judgment. Surgery must be paused to wait for the degradation to clear.
@@ 3.3 Data Statistics and Analysis
To facilitate fine-grained research, we further divided samples into multiple subsets based on different degradation levels and attributes. Specifically, for desmoking and defogging tasks, surgeons divide corresponding samples into four levels (from easy to difficult, i.e., Level 1-4) based on the area and degree of smoke or fog obstruction and the degree of interference with the operating zone. For desplashing, the samples are grouped into four subsets (i.e., $T_{blood}$, $T_{fat}$, $T_{bile}$, and $T_{fluid}$) based on the type of substances splashed onto the lens, including blood, fat, bile, and tissue fluid. We elaborate on the definition and classification of different levels of degradation in Table 2. In addition, Fig. 3 exhibits typical examples of different difficulty levels and each type of splash interference. Fig. 4 also illustrates the data statistics of SurgClean from two medical sites, showing the distribution of video sizes across surgical procedure categories and further subdivisions by specific procedure types. More sample exhibitions and information statistics of SurgClean dataset can be found in Supp. B.
Furthermore, we summarize and perform comparisons with other real-world ESIR datasets. As illustrated in Table 1, existing ESIR datasets focus solely on desmoking, while SurgClean comprehensively embraces multiple types of image degradation in real-world surgical procedures. In addition, datasets with paired reference labels provide more significant support for model fine-tuning. Compared to other datasets with a single data source, our data source comes from various types of endoscopic surgery, e.g., gallbladder, pancreatectomy, and esophageal, to accommodate data diversity. SurgClean comprises a variety of interference surgical scenarios to promote comprehensive evaluation. Additionally, the paired reference labels facilitate the training and evaluation of deep restoration models. We also perform a visual comparison of degraded images in our dataset with other datasets in Supp. C.
Figure 4: Data statistics of SurgClean from two medical sites, categorized by surgical procedure and detailed by surgical types.